OpenStudy (anonymous):

Which of the following is not a Pythagorean triple? A) 3 – 4 – 5 B)5 – 12 -13 C)7 – 24 – 25 D)8 – 15 – 18

OpenStudy (allank):

You can test each answer quickly: a^2+b^2=c^2

OpenStudy (allank):

And D seems to fit.

OpenStudy (anonymous):

A) 3^2 + 4^4 =? 5^5 9 + 16 = 25 This one is correct. B) 5^2 + 12^2 =? 13^2 25 + 144 = 169 This one is correct. C) 7^2 +24^2 =? 25^2 49 + 576 = 625 This one is correct. D) 8^2 + 15^2 =? 18^2 64 + 225 =X 324 This one is incorrect.

OpenStudy (anonymous):

@xVxNICKxVx so which one should i choose?

OpenStudy (anonymous):

You're looking for the one that does not conform to the a^2 + b^2 = c^2 formula. Which one of the equations was incorrect?

OpenStudy (anonymous):

D, thanks i get it

OpenStudy (anonymous):

:)